OpenStudy (anonymous):

given the terms a1 = 5 and a 4= 14 of an arithmetic sequence what is a20

OpenStudy (ghazi):

\[T _{20}=a _{1}+(n-1)d\] we need d

OpenStudy (ghazi):

now, \[a _{4}=a _{1}+(n-1)d, a _{4}=5+(4-1)d, 14=5+(3)d\] find d and you will have T 20 :)

OpenStudy (ghazi):

\[T _{20}=5+(20-1)d\]

OpenStudy (anonymous):

wait..... it is multiple choice it could be 62, 65,171, 185

OpenStudy (ghazi):

well calculate it :) your d=3 i guess then put d in the equation and find it :)

OpenStudy (anonymous):

wait what?? so t20=5+(19)3

OpenStudy (ghazi):

correct

OpenStudy (anonymous):

I got 62

OpenStudy (ghazi):

thats the answer :)

OpenStudy (anonymous):

awesome thank you!!! can you help me with more!!! i am so behind in my math class!

OpenStudy (ghazi):

keep posting :) don't follow up